# Exploring the evolution of red and blue galaxies in different cosmic web environments using IllustrisTNG simulation

### v1

@@ 4.1 Evolution of red and blue fractions in different cosmic web environments
In this subsection, we examine the evolution of the fraction of red and blue galaxies in various environments within the cosmic web. The red fraction (RF) and blue fraction (BF) in any cosmic web environment are respectively defined as $RF=\frac{n_{R}}{n_{R}+n_{B}}$ and $BF=\frac{n_{B}}{n_{R}+n_{B}}$. Here, $n_{R}$ and $n_{B}$ are the number of red and blue galaxies in the same environment. The top left and bottom left panels of Figure 3 illustrate these fractions across different cosmic web environments. Although the blue fraction can be obtained by simply subtracting the red fraction from one, they are separately shown here for the sake of completeness.
At redshift $z=3$, the blue fraction is highest in clusters, followed by filaments, sheets, and voids. Specifically, the blue fraction is $\sim 48\%$ in clusters and around $40\%$ in voids. Although the differences in blue fraction across environments at $z=3$ are not substantial, the trend suggests that overdense regions were more conducive to star formation in earlier times. This may arise due to the presence of larger gas reservoirs in these regions.
As redshift decreases, the blue fraction increases in all cosmic web environments. However, this trend reverses at lower redshifts, with the blue fraction declining and the red fraction rising. The blue fraction decreases and the red fraction increases at specific redshifts, with inflection points occurring at $z=1.5$ for clusters, $z=0.7$ for filaments and sheets, and $z=0.5$ for voids. At $z\leq 1.5$, red fractions dominate in clusters, followed by filaments, sheets, and voids. By $z=0$, the red fraction becomes $75\%$ in clusters, $45\%$ in filaments, $34\%$ in sheets, and $30\%$ in voids.
These results indicate an environment-dependent trend in galaxy colour transformation. Physical mechanisms that suppress star formation are more dominant in clusters and become effective earlier. This may be due to gradual gas depletion from ongoing star formation and faster gas exhaustion from interaction-induced starbursts in high-density regions. Suppression of star formation is less effective in filaments and least effective in sheets and voids, as evidenced by a high blue fraction ($\sim 68\%$) in these environments. This is also supported by a relatively smaller reduction in the blue fraction in sheets and voids from their respective peak values ($\sim 78\%$ at $z=0.5$).
@@
Figure 5: Same as Figure 4 but for blue fraction.
We now divide our galaxy sample at each redshift into two distinct mass bins: $9<\log(\frac{M_{*}}{M_{\odot}})\leq 10.5$ and $1210.5<\log(\frac{M_{*}}{M_{\odot}})<$ 12 to examine the evolution of red and blue fractions in lower and higher mass galaxies separately. Results for these two mass bins are presented in the middle and right panels of Figure 3. The middle panels display the evolution of red and blue fractions for the lower mass bin, which includes approximately $\sim 90\%$ galaxies of the entire sample (Table 2). The results for the lower mass bin are quite similar to those for the entire sample for this reason.
The results for the higher mass bin, containing about $\sim 10\%$ of the galaxies, are shown in the right panels of Figure 3. Even for the galaxies in the higher mass bin at $z=3$, we observe a higher blue fraction and a lower red fraction in clusters compared to filaments and sheets. However, the red fraction steadily increases across all environments as redshift decreases, indicating that star formation is gradually suppressed in massive galaxies regardless of their environment. Notably, the blue fraction in voids increases between $z=2$ and $z=1$, suggesting that massive galaxies in voids may follow different evolutionary trajectories compared to those in other cosmic web environments.
@@
We also examine the red and blue fractions as functions of stellar mass within each geometric environment. Figure 4 and Figure 5 show these results across different redshifts. The bottom right panels of these figures reveal that at $z=3$, clusters generally have a higher blue fraction and a lower red fraction across nearly all masses. The blue fraction in filaments is lower than in clusters but higher than in sheets and voids throughout the entire mass range. However, the larger error bars for voids and sheets limit the statistical significance of these differences. Between $z=3-1.5$, we observe that in each geometric environment, the blue fraction increases significantly in lower mass galaxies ($\log(\frac{M_{*}}{M_{\odot}}\leq 10.5$). The blue fraction in lower mass galaxies do not exhibit a strong dependence on the geometric environment during this period. A clear environment-dependent trend emerges at $z<1$. We find that in clusters, the red fraction of lower mass galaxies increases from approximately $\sim 40\%$ to $\sim 70\%$ between redshifts 1 and 0. For higher mass galaxies, the red fraction rises from about $\sim 80\%$ to $\sim 95\%$ in all environments over the same period. Further, $\sim 40\%$ of galaxies in filaments are red, compared to around $\sim 30\%$ in sheets and voids. This indicates that the quenching of lower mass galaxies is strongly influenced by their environment at $z<1$, while quenching in more massive galaxies is less affected by their embedding geometric environments during the same period.
@@ 4.3 Evolution of relative red and blue fractions in different cosmic web environments
The red and blue fractions provide the proportion of two populations in any particular environment. We can use the relative red fraction (RRF) and the relative blue fraction (RBF) to study the relative abundance of red and blue galaxies in various cosmic web environments. The RRF and RBF are respectively defined as $RRF=\frac{(n_{R})_{i}}{\sum_{i=1}^{4}(n_{R})_{i}}$ and $RBF=\frac{(n_{B})_{i}}{\sum_{i=1}^{4}(n_{B})_{i}}$, where $(n_{R})_{i}$ and $(n_{B})_{i}$ are the numbers of red and blue galaxies in $i^{th}$-type environment. Here $i$ runs from $1$ to $4$ covering 4-types of cosmic web environment. We examine the evolution of the RRF and RBF across different cosmic web environments, as shown in the top and bottom left panels of Figure 6. The RRF and RBF represent the proportions of red and blue galaxies, respectively, within a specific environment compared to other cosmic environments. The top and bottom left panels of Figure 6 reveal that in filaments and sheets, the RRF decreases and the RBF increases with decreasing redshift. Conversely, clusters show an opposite trend where the RRF increases and the RBF decreases as redshift decreases. Throughout the entire redshift range, filaments consistently have the highest RBF. At lower redshifts ($z<1$), clusters exhibit the highest RRF compared to other environments.
@@
Figure 7 illustrates the evolution of the relative fractions of galaxies in filaments, sheets, clusters, and voids for the entire sample and the two distinct mass bins. The left panel of Figure 7 shows that filaments generally have the highest galaxy abundance, followed by clusters, sheets, and voids across the entire redshift range. The proportion of galaxies in filaments and sheets increases with decreasing redshift, while the proportion in clusters decreases. These trends are also evident in the lower mass bin, as depicted in the middle panel of Figure 7. In the high mass bin, the proportion of galaxies in filaments and clusters evolves differently, as shown in the right panel of Figure 7. There is a crossover around $z\sim 1$, after which filaments have a higher proportion of galaxies compared to clusters. Prior to $z=1$, clusters dominate in terms of galaxy proportion.
@@
[…]
The anomalous results in the right panels of Figure 6 are explained by the trends observed in the right panel of Figure 7 and Figure 3. For high mass galaxies, RF increases and BF decreases in all types of cosmic web environments, with comparable values across different environments (Figure 3). Consequently, the magnitudes of the RRF and RBF are primarily determined by the proportions of galaxies in different environments. The trends observed in the right panel of Figure 7 are reflected in the top and bottom right panels of Figure 6, with crossovers in RRF and RBF occurring at different redshifts due to the varying red and blue fractions in filaments and clusters.
@@
We now examine the RRF and RBF across different cosmic web environments as functions of stellar mass and their evolution from redshift $3$ to 00. The various panels in Figure 8 display the RRF as a function of stellar mass in different cosmic web environments at different redshifts. The results indicate that in both filaments and sheets, the RRF is lower at higher masses and higher at lower masses. The RRF in filaments and sheets at lower masses decreases with decreasing redshift. Conversely, clusters exhibit an opposite trend, where the RRF is higher at higher masses and lower at lower masses. In clusters, the RRF at higher masses remains relatively stable with redshift, while at lower masses, it gradually increases as redshift decreases. At $z\sim 1$, the RRF at lower masses becomes comparable between filaments and clusters. However, below this redshift, clusters tend to dominate at lower masses, while filaments become more prominent at intermediate masses. At low redshifts, the RRF in clusters exceeds that of all other environments at both low and high masses. This suggests that low-mass red galaxies in clusters may represent quenched satellites, while high-mass red galaxies may be massive centrals. Interestingly, the RRF in clusters and filaments shows two distinct crossovers at intermediate masses. Between these crossovers ($1110.5<\log(\frac{M_{*}}{M_{\odot}})<$ 11), the RRF in filaments surpasses that of all other environments. It may be worthwhile to mention here that the right panels of Figure 6 indicate that, at lower redshifts, both RRF and RBF dominate in filaments for masses $\log(\frac{M_{*}}{M_{\odot}})>10.5$. However, a closer examination reveals that this dominance is confined to the mass range $1110.5<\log(\frac{M_{*}}{M_{\odot}})<$ 11. There are only a small number of red and blue galaxies with masses $\log(\frac{M_{*}}{M_{\odot}})>11$, and these primarily reside in clusters.
@@
We present the RBF as a function of stellar mass across different cosmic web environments and redshifts in the various panels of Figure 9. Our analysis reveals that in both filaments and sheets, the RBF is lower at higher masses and higher at lower masses. Specifically, the RBF in filaments and sheets at lower masses increases as redshift decreases. Between redshift $z=1-0$, filaments exhibit a significantly higher RBF ($\sim 0.55$) compared to other environments for galaxies having $\log(\frac{M_{*}}{M_{\odot}})<11$. This indicates that filaments host a diverse population of galaxies, encompassing both passive and star-forming galaxies. The less aggressive environmental quenching in filaments allows for a wider range of evolutionary stages. Interestingly, despite being the densest environment, clusters show the highest RBF for high-mass galaxies ($\log(\frac{M_{*}}{M_{\odot}}>11$) even at lower redshifts. This suggests that clusters are highly dynamic environments where interactions or mergers might temporarily reignite star formation activity in some high mass galaxies.
Figure 10: This figure shows the PDF of $(u-r)$ colour in different cosmic web environments at different redshifts.
@@ 4.5 Evolution of colour across different cosmic web environments
We present the distributions of $(u-r)$ colour across various cosmic web environments at different redshifts, ranging from $z=3$ to $z=0$, as shown in different panels of Figure 10. At $z=3$, the $(u-r)$ colour distribution in each environment is unimodal, with a peak around $(u-r)\sim 0.7$. As redshift decreases, this peak shifts towards higher $(u-r)$ values, indicating the aging of stellar populations. By $z=2$, the colour distribution begins to show bimodality. The bimodality becomes most pronounced in clusters, followed by filaments, sheets, and voids. In clusters, the peak corresponding to the red sequence grows rapidly, eventually surpassing the peak of the blue cloud in this environment at $z=0.5$. The emergence of distinct bimodality across all environments suggests that the colour transformation can occur in any cosmic web environment.
In the top left panel of Figure 11, we show the median $(u-r)$ colour of galaxies in various cosmic web environments as a function of redshift. Since $z\sim 2$, the median colour in clusters begins to diverge from that in other environments and increases rapidly after $z>1$. Meanwhile, the median colour in filaments starts to differ from that in sheets and voids after redshift 1. At $z<1$, clusters exhibit the highest median colour, followed by filaments, sheets, and voids.
The top right and bottom panels of Figure 11 illustrate the evolution of the median colour in each environment for lower and higher mass bins. The lower mass bin shows trends similar to the combined sample due to the large number of galaxies in this bin. The bottom panel reveals that the median colour for more massive galaxies ($1210.5<\log(\frac{M_{*}}{M_{\odot}})<$ 12) increases steadily at $z<1$, with little variation across different environments. This suggests that the colour evolution of the more massive galaxies is primarily driven by mass and is less dependent on their environment. This reaffirms our findings in subsection 4.1 and subsection 4.2.
@@ 4.6 Evolution of stellar mass across diverse cosmic web environments
In the top left panel of Figure 13, we show the evolution of the median stellar mass of galaxies across different cosmic web environments. It shows that as redshift decreases, the median mass of galaxies increases in all environments. At each redshift, the median mass is highest in clusters, followed by filaments, sheets, and voids. The top right and bottom panels of Figure 13 show the results for lower and higher mass bins, respectively. In the top right panel, we observe that the median mass of galaxies in filaments and sheets increases steadily between $z=1$ and $z=0$. However, at $z<1$, the growth rate of the median mass for lower mass galaxies in clusters slows down significantly compared to those in filaments and sheets. The bottom panel shows that the median stellar mass for high mass galaxies remains stable between $z=3$ and $z=1.5$, followed by a gradual increase up to the present. This suggests that these galaxies undergo a period of minimal mass growth during the earlier stages of cosmic evolution, likely indicating that major mergers are less frequent at higher redshifts. The steady increase in median stellar mass at $z<1.5$ points to more significant mass accumulation in later stages, possibly due to mergers and accretion.

### v2

@@ 4.1 Evolution of red and blue fractions in different cosmic web environments
In this subsection, we examine the evolution of the fraction of red and blue galaxies in various environments within the cosmic web. The top left and bottom left panels of Figure 5 illustrate the red and blue fractions (RF and BF) across different cosmic web environments. Although the blue fraction can be obtained by simply subtracting the red fraction from one, they are separately shown here for the sake of completeness.
From the bottom left panel of Figure 5, we observe that at redshift $z=3$, the blue fraction is highest in clusters, followed by filaments, sheets, and voids. Specifically, the blue fraction is $\sim 48\%$ in clusters and around $35\%$ in voids. Although the differences in blue fraction across environments at $z=3$ are not substantial, the trend suggests that overdense regions were more conducive to star formation in earlier times. This may arise due to the presence of larger gas reservoirs in these regions.
The first column of Figure 5 shows that as redshift decreases, the blue fraction increases in all cosmic web environments. However, this trend reverses at lower redshifts, with the blue fraction declining and the red fraction rising. The blue fraction decreases and the red fraction increases at specific redshifts, with inflection points occurring at $z=1.5$ for clusters, $z=0.7$ for filaments and sheets, and $z=0.5$ for voids. At $z\leq 1.5$, red fractions dominate in clusters, followed by filaments, sheets, and voids. By $z=0$, the red fraction becomes $75\%$ in clusters, $45\%$ in filaments, $34\%$ in sheets, and $30\%$ in voids.
These results indicate an environment-dependent trend in galaxy colour transformation. At $z=3$, clusters exhibit the highest blue galaxy fraction, followed by filaments, sheets, and voids. This trend suggests that early clusters may have hosted more star-forming galaxies, potentially due to higher gas availability. The blue fraction increases across all environments until redshift $z=1.5$, which likely corresponds to the peak of cosmic star formation activity. After $z=1.5$, the blue fraction begins to decline, with red galaxies becoming dominant in clusters. This shift may reflect the onset of processes that suppress star formation, such as gas depletion, AGN feedback, or galaxy interactions, which are expected to be more prevalent in dense environments. The timing of this transition varies across environments, with filaments and sheets showing later changes compared to clusters. This pattern suggests that the influence of environment on galaxy evolution may depend on local density, although we caution that our analysis does not establish causality. The blue fraction remains highest in voids until $z<0.5$, after which an increase in red galaxy fraction is observed. These trends are consistent with the idea that quenching processes become effective earlier in denser environments. The inflection points seen in Figure 5 thus reflect a sequence in galaxy evolution that aligns with increasing environmental density, though we emphasize that these are observed trends and not direct evidence of specific physical mechanisms.
@@
Figure 7: Same as Figure 6 but for red fraction.
We now divide our galaxy sample at each redshift into two distinct mass bins: $9<\log(\frac{M_{*}}{M_{\odot}})\leq 10.5$ and $10.5<\log(\frac{M_{*}}{M_{\odot}})<12$ to examine the evolution of red and blue fractions in lower and higher mass galaxies separately. Results for these two mass bins are presented in the second and third columns of Figure 5. The second column of Figure 5 displays the evolution of red and blue fractions for the lower mass bin, which includes approximately $\sim 90\%$ galaxies of the entire sample (Table 2). For this reason, the results for the lower-mass bin closely resemble those of the full sample.
The results for the higher mass bin, containing about $\sim 10\%$ of the galaxies, are shown in the third column of Figure 5. The third column of Figure 5 shows that even for the galaxies in the higher mass bin at $z=3$, we observe a higher blue fraction and a lower red fraction in clusters compared to filaments and sheets. However, the red fraction steadily increases across all environments as redshift decreases, indicating that star formation is gradually suppressed in massive galaxies regardless of their environment. Interestingly, we note that massive galaxies in voids experience an increase in star formation between $z=2$ and $z=1$, unlike similarly massive galaxies in other environments. The sharp increase also indicates that massive galaxies in voids may follow different evolutionary trajectories compared to those in other cosmic web environments. The gas from the expanding void regions can significantly enhance star formation activity in galaxies residing within them [126, 127]. The absence of the $z=3$ point in the third column of Figure 5 is due to the lack of galaxies with masses between $10.5<\log(\frac{M_{*}}{M_{\odot}})<12$ in voids at $z=3$.
@@
We analyze the red and blue galaxy fractions as functions of stellar mass across different cosmic web environments, as shown in Figures 6 and 7. At $z=3$, clusters exhibit a higher blue fraction and lower red fraction across most mass ranges. Filaments show intermediate blue fractions that are lower than clusters but higher than sheets and voids. However, the larger uncertainties due to a smaller number of galaxies in sheets and voids reduce the statistical significance of these differences.
From $z=3$ to $z=1.5$, the blue fraction rises significantly for low-mass galaxies ($\log(M_{\star}/M_{\odot})\leq 10.5$) in all environments, with minimal environmental dependence. A clear environment-dependent trend emerges at $z<1$ for low-mass galaxies, where the red fraction increases in clusters from $\sim 40\%$ at $z=1$ to $\sim 70\%$ at $z=0$. At $z=0$, $\sim 40\%$ low mass galaxies in filaments are red compared to $\sim 30\%$ in sheets and voids. In contrast, for high-mass galaxies ($\log(M_{\star}/M_{\odot})>10.5$), the red fraction increases across all environments from $\sim 80\%$ to $\sim 95\%$ suggesting that quenching in massive galaxies is largely independent of environment.
We observe a peak in the blue fraction at $\log(M_{\star}/M_{\odot})\sim 10.5$ between $z=2$ and $z=1$ across all cosmic web environments (Figure 6). This indicates that intermediate-mass galaxies were the most actively star-forming population during this epoch, regardless of their location within the cosmic web. The presence of this peak suggests that galaxies in this mass range had sufficient gas reservoirs and moderate feedback regulation, enabling sustained star formation before the onset of widespread quenching mechanisms at lower redshifts. The sharp increase in the red fraction at $\log(M_{\star}/M_{\odot})\sim 10.5$ (Figure 7) marks a critical transition in galaxy evolution, where galaxies shift from being primarily star-forming to increasingly quenched. Below this mass threshold, environmental factors dominate quenching, while internal mechanisms like mass quenching and AGN feedback take over above it [7, 8, 128, 129, 130, 131].
@@ 4.3 Evolution of relative red and blue fractions in different cosmic web environments
We examine the evolution of the RRF and RBF across different cosmic web environments, as shown in the top and bottom left panels of Figure 8. The top and bottom left panels of Figure 8 reveal that in filaments and sheets, the RRF decreases and the RBF increases with decreasing redshift. Conversely, clusters show an opposite trend where the RRF increases and the RBF decreases as redshift decreases. Throughout the entire redshift range, filaments consistently have the highest RBF. At lower redshifts ($z<1$), clusters exhibit the highest RRF compared to other environments.
@@
The anomalous results in the right panels of Figure 8 can be explained by the trends observed in the right panel of Figure 3 and Figure 5. For high mass galaxies, RF increases and BF decreases in all types of cosmic web environments, with comparable values across different environments (Figure 5). Consequently, the magnitudes of the RRF and RBF are primarily determined by the proportions of galaxies in different environments. The trends observed in the right panel of Figure 3 are reflected in the top and bottom right panels of Figure 8, with crossovers in RRF and RBF occurring at different redshifts due to the varying red and blue fractions in filaments and clusters.
@@
We examine the evolution of the RRF and RBF as functions of stellar mass and cosmic web environment from $z=3$ to $z=0$ (Figure 9, Figure 10). In filaments and sheets, the RRF is relatively higher at low stellar masses compared to higher masses and decreases with decreasing redshift, indicating that low-mass galaxies in these environments tend to remain blue for longer periods. Conversely, in clusters, the RRF is initially higher at high masses but increases at low masses over time, suggesting that environmental quenching of low-mass galaxies becomes more prominent at late times.
At $z\sim 1$, the RRF at low masses becomes comparable in filaments and clusters. By $z=0$, however, clusters dominate the red population at both low and high stellar masses, while filaments show a peak in RRF at intermediate masses ($10.5<\log(M_{\star}/M_{\odot})<11$). Two notable crossovers in the RRF curves at $z=0$ illustrate this shift: at low masses, clusters overtake filaments in RRF, indicating intensified quenching in dense environments; and at intermediate masses, filaments surpass clusters, emerging as the primary hosts of quiescent galaxies in that mass range.
RBF trends provide a complementary view. Clusters consistently show lower RBF values at low stellar masses across redshifts, which is consistent with clusters hosting lower fractions of blue, low-mass galaxies. In contrast, filaments show the highest RBF at intermediate and low masses for $z<1$, highlighting their transitional role hosting both star-forming and quenched galaxies. This contrasts with clusters, which are associated with efficient quenching and consistently exhibit low RBF values compared to filaments across redshifts. Voids, on the other hand, contain a very small number of galaxies overall, resulting in low RRF and RBF values at all stellar masses and redshifts. While their low-density nature suggests a star-forming preference, the small sample size limits statistical interpretation.
These trends support the idea that mass-driven and environment-driven quenching processes operate in tandem, but on different timescales and with varying efficiencies across environments. Filaments appear to quench galaxies earlier, particularly at low stellar masses, while clusters become increasingly effective at quenching low-mass galaxies at later times. For intermediate-mass galaxies, filaments play a dominant role significantly influencing their evolutionary pathways. Our results are consistent with previous findings [132, 130], which show that low-mass galaxies in dense environments quench rapidly, whereas intermediate-mass galaxies are more sensitive to the surrounding large-scale structure. Interestingly, clusters also exhibit the highest RBF at the high-mass end ($\log(M_{\star}/M_{\odot})>11$) even at low redshift, possibly reflecting episodes of renewed star formation triggered by mergers or dynamical interactions in these dense environments.
@@
Figure 11: This figure shows the PDF of (ur)(u-r) colour in different cosmic web environments at different redshifts.
@@
The trends observed in all the galaxy fractions across cosmic web environments may, in part, be driven by variations in local density, since denser regions naturally host different types of galaxies compared to underdense areas. To disentangle the effects of local density from those of the geometric environment, it is essential to examine the fractions F, RF, BF, RRF, and RBF within narrow density bins. This approach would allow us to assess whether the observed environmental trends are intrinsic to the cosmic web environment or simply a reflection of underlying density variations, providing a clearer picture of the role played by cosmic web in galaxy evolution. We address this issue in detail in subsection A.1 by analyzing the galaxy fractions within narrow density bins to isolate the influence of cosmic web geometry from that of density.
@@ 4.5 Evolution of colour across different cosmic web environments
We present the distributions of (ur)(u-r) colour across various cosmic web environments at different redshifts, ranging from $z=3$ to $z=0$, as shown in different panels of Figure 11. At $z=3$, the (ur)(u-r) colour distribution in each environment is unimodal, with a peak around $(u-r)\sim 0.7$. As redshift decreases, this peak shifts towards higher (ur)(u-r) values, indicating the aging of stellar populations. By $z=2$, the colour distribution begins to show bimodality. The bimodality becomes most pronounced in clusters, followed by filaments, sheets, and voids. In clusters, the peak corresponding to the red sequence grows rapidly, eventually surpassing the peak of the blue cloud in this environment at $z=0.5$. The emergence of distinct bimodality across all environments suggests that the colour transformation can occur in any cosmic web environment.
In the top left panel of Figure 12, we show the median (ur)(u-r) colour of galaxies in various cosmic web environments as a function of redshift. Since $z\sim 2$, the median colour in clusters begins to diverge from that in other environments and increases rapidly after $z>1$. Meanwhile, the median colour in filaments starts to differ from that in sheets and voids after redshift 1. At $z<1$, clusters exhibit the highest median colour, followed by filaments, sheets, and voids.
The top right and bottom panels of Figure 12 illustrate the evolution of the median colour in each environment for lower and higher mass bins. The lower mass bin shows trends similar to the combined sample due to the large number of galaxies in this bin. The bottom panel reveals that the median colour for more massive galaxies ($10.5<\log(\frac{M_{*}}{M_{\odot}})<12$) increases steadily at $z<1$, with little variation across different environments. This suggests that the colour evolution of the more massive galaxies is primarily driven by mass and is less dependent on their environment. This reaffirms our findings in subsection 4.1 and subsection 4.2.
@@ 4.6 Evolution of stellar mass across diverse cosmic web environments
In the top left panel of Figure 14, we show the evolution of the median stellar mass of galaxies across different cosmic web environments. It shows that as redshift decreases, the median mass of galaxies increases in all environments. At each redshift, the median mass is highest in clusters, followed by filaments, sheets, and voids. The top right and bottom panels of Figure 14 show the results for lower and higher mass bins, respectively. In the top right panel, we observe that the median mass of galaxies in filaments and sheets increases steadily between $z=3$ to $z=1$. However, at $z<1$, the growth rate of the median mass for lower mass galaxies in clusters slows down significantly compared to those in filaments and sheets. The bottom panel shows that the median stellar mass for high mass galaxies remains stable between $z=3$ and $z=1.5$, followed by a gradual increase up to the present. This suggests that these galaxies undergo a period of minimal mass growth during the earlier stages of cosmic evolution, likely indicating that major mergers are less frequent at higher redshifts. The steady increase in median stellar mass at $z<1.5$ points to more significant mass accumulation in later stages, possibly due to mergers and accretion.